OpenStudy (anonymous):

Which fraction is equivalent to 6/9? A.12/18 B.3/4 C.1/3 D.18/36

OpenStudy (anonymous):

A

OpenStudy (anonymous):

what is 18/30 in lowest terms?

OpenStudy (anonymous):

3/5

OpenStudy (anonymous):

divide both numerator and denominator by 6

OpenStudy (anonymous):

lol your too smart

OpenStudy (anonymous):

what grade are you?

OpenStudy (anonymous):

im in alg 2 so this is easy for me

OpenStudy (anonymous):

though compaired to calculus, alg 2 is easy as pellet

OpenStudy (anonymous):

What value of a makes the equation true? 30/45 =6/a

OpenStudy (anonymous):

this has autocorrect?????? 0.0 i said pellet

OpenStudy (anonymous):

wtf man 0.0

OpenStudy (anonymous):

S H I T

OpenStudy (anonymous):

there we go :)

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

do it, type S H I T normally in lowercase, youll see

OpenStudy (anonymous):

its spells pellet

OpenStudy (anonymous):

idk why but it does :|

OpenStudy (anonymous):

pellet

OpenStudy (anonymous):

XD

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

doooo it XD

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

:o

OpenStudy (anonymous):

you didnt do it

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

What value of a makes the equation true? 30/45 =6/a

OpenStudy (anonymous):

do what lol

OpenStudy (anonymous):

6/9 which is also simplified to 2/3

OpenStudy (anonymous):

pellet

OpenStudy (anonymous):

type S H I T normally

OpenStudy (anonymous):

into chat

OpenStudy (anonymous):

yea i did lol

OpenStudy (anonymous):

it turns into pellet

OpenStudy (anonymous):

pellet

OpenStudy (anonymous):

pellet pellet pellet

OpenStudy (anonymous):

lmao XD

OpenStudy (anonymous):

im such a dork XD

OpenStudy (anonymous):

i sent the answer

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

What value of a makes the equation true? 30/45 =6/a

OpenStudy (anonymous):

yeah 6/9 or 2/3

OpenStudy (anonymous):

its not an answer

OpenStudy (anonymous):

It's 12/18

OpenStudy (anonymous):

a9 b6 c30 d225

OpenStudy (anonymous):

i said it was A at the top of the chat lol, first message

OpenStudy (anonymous):

its A

OpenStudy (anonymous):

6/9

OpenStudy (anonymous):

9 is a

OpenStudy (anonymous):

What value of a makes the equation true? 30/45 =6/a

OpenStudy (anonymous):

becaus eit says 6/a and if the answer is 6/9 then a is 9

OpenStudy (anonymous):

YOUR ANSWER IS A.9

OpenStudy (anonymous):

what is 5/9 x 3/10?

OpenStudy (anonymous):

i dont multiply them well lmao XD hold on

OpenStudy (anonymous):

i just multiplied them straight across and got 15/90 which simplifies to 3/18 which again simplifies to 1/6

OpenStudy (anonymous):

k lmao

OpenStudy (anonymous):

idk why but when i devided 90 by 15 it didnt work at first 0.0 so i just did a bunch of extra work XD

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

is that all your questions?

OpenStudy (anonymous):

hold up i got the answer :)

OpenStudy (anonymous):

its 1/6, is it not?

OpenStudy (anonymous):

please educate me, im fluttering horrible at multiplying fractions XD

OpenStudy (anonymous):

:|

OpenStudy (anonymous):

i put flutterING

OpenStudy (anonymous):

:|

OpenStudy (anonymous):

lol its 1/6

OpenStudy (anonymous):

F U C K I N G

OpenStudy (anonymous):

man...this site makes me crack up XD

OpenStudy (anonymous):

ok good lol

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

5/6 divided by 1/3

OpenStudy (anonymous):

:o

OpenStudy (anonymous):

5/2

OpenStudy (anonymous):

i think

OpenStudy (anonymous):

because you take it and flip the fraction over (just one) and multiply it then simplify

OpenStudy (anonymous):

its 2 1/2

OpenStudy (anonymous):

k now we working with LCD

OpenStudy (anonymous):

ready :)

OpenStudy (anonymous):

well that makes sense, you didnt say whether it to be a mixed number or an improper fraction, 5/2 is equivalent to 2 1/2

OpenStudy (anonymous):

i suppose as ready as ill ever be

OpenStudy (anonymous):

XD

OpenStudy (anonymous):

what is the LCD of 3/4 and 7/10

OpenStudy (anonymous):

2

OpenStudy (anonymous):

XD my bad forgot to tell you

OpenStudy (anonymous):

forgot to tell me what?

OpenStudy (anonymous):

a10 b 80 c20 d40

OpenStudy (anonymous):

20

OpenStudy (anonymous):

4 cannot go into 10 and the next number 10 can go into is 20, and 4x5 = 20 , so its 20

OpenStudy (anonymous):

k next question is adding which i got hen\

OpenStudy (anonymous):

then we will be on last question

OpenStudy (anonymous):

ok, well ill brb. i have to unload wood from the truck real quick

OpenStudy (anonymous):

k

OpenStudy (anonymous):

Dude we got a 88% :D

OpenStudy (anonymous):

yay! :)